Calculate the molar concentrations of oxalic acid (COOH)2, HOOCCO2-, (CO2)22-, H3O+, and OH- in 0.15 M (COOH)2 (aq).

This is a special case of general diprotic acid dissociation. First, we need to find the pKa1 and pKa2 in the water at room temperature, which is available online e.g. Wikipedia (oxalic acid). pKa1 = 1.23, pKa2 = 4.19.

Either solve numerically or analytically. Report here [H3O+] = ______ M. 2 sig. figures.
Answer #1

0.15m o Nauj K9 Dissociation of oxalic acid ? tuo H2C2O4 Kaul HC2O4 +HT Initial Final (Equilibrium) 0-15- * XT X be the amounGiven, POSH РКау - 1.23. We know rocha Қа, а M20.00 Pka, = -log(kas) — дана | 23 с - log (Kal) 0+058. І. • 058 - д45-X. х agKaz HC 204 Initial 0.0685m Eqm. 0.0685- 685 %. C2042 +ht. Dell go od mo CY 10%) pa y y be the degree of dissociation PKaza 4.5.13x10-12. PH= -log [t zot 2.716 -log [H 3 ot [H3O+] = 1.95*10-3 [OH-] = KW [ht 10-14 1.95810-3